Grade 11Mechanics

two trains each of length 100 metres are running on parallel tracks. One overtakes the other in 20seconds and one crosses the other in 10 secondss. Calculate the speed of both trains.

Harsh Raj
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
The right answer is 450 m 
Here is my clarification. 
Utilizing idea of relative speed, 
u = 0 (both the trains have same speed); S = x(i.e. unique separation between the trains) + 2*400(the prepare B will likewise venture to every part of the length of its own and furthermore the length of the prepare A) 
Utilizing second condition of movement, 
S = ut + 1/2 at^2 
=> x + 800 = 1/2 * 1 * 2500 = 1250 m 
=> x = 450 m 
In this manner the first separation between the trains ought to be 450 m.